For a recent 2-year period, the balance sheet of Skysong Company showed the following stockholders' equity data at December 31 (in millions). 


                                           2020  2019

 Additional paid-in capital      $930 $843

 Common stock                    651 642

 Retained earnings                7,210 5,220

 Treasury stock                     1,850 945

 Total stockholders' equity     $6,941 $5,760


 Common stock shares issued  217 214

 

 Common stock shares authorized 500 500

 Treasury stock shares               37 27

 

 (a) Answer the following questions.

 (1) What is the par value of the common stock? (Round par value to 2 decimal places, e.g. $3.15)

 Par value of common stock



(2) What is the cost per share of treasury stock at December 31, 2020, and at December 31, 2019? 

(b) Prepare the stockholders' equity section at December 31, 2020. (Enter account name only and do not provide descriptive information.) 
Answer #1

a1) Par value of common Stock = 651/217 = 3

a2) Cost per share of treasury stock

December 31,2020 December 31,2019
Cost per share 1850/37 = 50 945/27 = 35

b)

Stockholder's equity
Paid in capital
Common Stock 651
Additional paid in capital 930
Total paid in capital 1581
Retained earnings 7210
Total paid in capital and retained earnings 8791
Less: Treasury stock -1850
Total Stockholders equity 6941